# Propaganda is All You Need

### v1

@@ Abstract
As ML is still a (relatively) recent field of study, especially outside the realm of abstract mathematics, few works have been conducted on the political aspect of LLMs, and more particularly about the alignment process and its political dimension. This process can be as simple as prompt engineering but is also very complex and can affect completely unrelated questions. For example, politically directed alignment has a very strong impact on an LLM’s embedding space and the relative position of political notions in such a space. Using special tools to evaluate general political bias and analyze the effects of alignment, we can gather new data to understand its causes and possible consequences on society. Indeed, by taking a socio-political approach, we can hypothesize that most big LLMs are aligned with what Marxist philosophy calls the ’dominant ideology.’ As AI’s role in political decision-making—at the citizen’s scale but also in government agencies—such biases can have huge effects on societal change, either by creating new and insidious pathways for societal uniformity or by allowing disguised extremist views to gain traction among the people.
@@
Paul Kronlund-Drouault École Normale Supérieure de Lyon paul.kronlund-drouault@ens-lyon.fr
@@ 1 Introduction
How many times, since the start of the current Artificial Intelligence (AI) era, have I been told by friends, professors, or even colleagues to ask a Large Language Model (LLM) to get an unbiased opinion about a subject? Indeed, in the collective doxa, the belief that LLMs are still ruled by deterministic and logical rules implies, a fortiori, that they must be neutral, or even ’apolitical,’ and that they thus represent a way to obtain an ’objective opinion’  [^1] Indeed, we’ll see in §5 how the nature of LLMs makes them a hard-to-define subject for critical thinking. While this might seem unimportant at first, it convinced me to dig a bit deeper into the underlying structure defining an LLM’s ideology and how it can be expressed through its output. In order to do this, we will also focus on a process called alignment (§3). Building upon multiple works highlighting the political nature of AI and pointing out specific biases, the goal of this research work is to understand the different ideological biases present in LLMs through a deep analysis of the inner workings of LLMs and particularly the alignment process, and to evaluate our findings using a socio-political framework to highlight their societal causes and possible consequences. We are first going to review the previous works on the matter, their approaches to bias analysis and political classification, and try to elaborate new methods to better understand the nature of ideological tendencies in LLMs. We are going to dive into the embedding space and other crucial AI concepts to study how political biases form and how they might be stored, both in the MLP (3Blue1Brown, 2024) and the model’s embedding space (Tennenholtz et al., 2024). Then, using both experimental data we gathered for this project and the findings of other papers on the subject, we are going to try and formulate a socio-political hypothesis to explain such biases, the need for and dangers of alignment, and the possible societal impact of political AI.
@@ 3.1 Unsupervised Alignment
As of right now, we don’t possess enough data to precisely measure the profound impact of particular data on a model aligned with unsupervised training since the effects of such training are hard to measure. Some early results we obtained for this work show that such training has an impact on a model’s embedding space, which basically contains the meanings the AI associates with the words we give it. In our tests, an LLM trained on Trotskyist data will have notions like "communism" and "Stalin" farther apart than, for example, a model trained on Marxist-Leninist data  [^4] Along this paper, we will use a lot of alignment references from different types of Marxist-aligned LLMs. The reason for this choice is the abundance of good-quality Marxist data available for free. (mar, ). Training also affects the space in ways we cannot yet understand, changing the position of sometimes completely unrelated notions.
@@ 3.2 ORPO / DPO Alignment
This concept of alignment has a lot of similarities with human-directed political influence. Some techniques, such as DPO (Rafailov et al., 2024) or ORPO (Jiwoo Hong and Thorne, 2023), use data sorted into ideological categories and assign weights to them. These techniques aim to modify the AI tendencies to align with them and can evoke quite literally propaganda posters from WWII or the Cold War, mixing desired ideologies with positive signals and the ’enemy’ position with negative brain stimuli. Most big players in today’s LLM market use such techniques to prevent behaviors they find harmful. This method is very powerful as it doesn’t require a lot of training to be effective, but I hypothesize that their impact is more on the ’surface level,’ acting like a roleplay prompt rather than completely changing the inner structure of the model.
As for the unsupervised alignment, we also measure a difference in embedding space, which, although smaller, affects a reduced set of vectors in more predictable ways.
@@ 4.1.2 Absolute Position
In order to obtain more traditional conservative-liberal-Marxist data on an LM’s biases, we use a different method that will provide less precise insights but will be grounded in real-world data.
@@ 5 Socio-Political Impacts
In the current state of AI, all major models are produced not by governments but by privately owned corporations. As with social networks, we can hypothesize that this condition will contribute to the *power shift* from governments to private entities by giving them much more influence. To prevent this, AI regulation would need not only to control an LLM’s output but also precisely measure its biases, and raise awareness around these issues. Indeed, in a lot of countries, political discourse—especially in election times—is heavily regulated (CC, ), and as AI can convey ideological messages, it might need to be subjected to the same types of rules.
@@ 5.1 A Marxian Analysis
Our analysis can even be deepened by bringing in the idea of hegemony, developed by Antonio Gramsci Gramsci (1971). Indeed, Language Models represent in the current context an essences of the hegemonic ideology of the ruling class. Through the complex tensors of the Multi-layer perceptron, the main points of the capitalist thought framework are encoded, with alignment and/or natural bias. Furthermore, Gramsci theorizes that class hegemony goes through a socio-cultural process that alters ideas themselves. This concept has many similarities with the embedding space tuning in the alignment process: An LLM’s world model is changed not only in its factual aspects but directly in its analysis pathways.  [^8] Further research should be conducted regarding the specific implications of artificial embedding space alterations on a model’s analysis. In the meantime, given the mathematical model of LLMs, we can already assume that they are correlated. See 3.1 for our (scarce) findings around this..

### v2

@@ Abstract
As Machine Learning (ML) is still a recent field of study, especially outside the realm of abstract Mathematics and Computer Science, few works have been conducted on the political aspect of large Language Models (LLMs), and more particularly about the alignment process and its political dimension. This process can be as simple as prompt engineering but is also very complex and can affect completely unrelated notions. For example, politically directed alignment has a very strong impact on an LLM’s embedding space and the relative position of political notions in such a space. Using special tools to evaluate general political bias and analyze the effects of alignment, we can gather new data to understand its causes and possible consequences on society. Indeed, by taking a socio-political approach, we can hypothesize that most big LLMs are aligned with what Marxist philosophy calls the ’dominant ideology.’ As AI’s role in political decision-making—at the citizen’s scale but also in government agencies—such biases can have huge effects on societal change, either by creating new and insidious pathways for societal uniformity or by allowing disguised extremist views to gain traction among the people.
@@
Paul Kronlund-Drouault pkd.unsuspicious.org
@@ 1 Introduction
How many times, since the start of the current Artificial Intelligence (AI) era, have I been told by friends, professors, or even colleagues to ask a Large Language Model (LLM) to get an unbiased opinion about a subject ? Indeed, in the collective doxa, the belief that LLMs are still ruled by deterministic and logical rules implies, a fortiori, that they must be neutral, or even ’apolitical,’ and that they thus represent a way to obtain an ’objective opinion’  [^1] Indeed, we’ll see in §5 how the nature of LLMs makes them a hard-to-define subject for critical thinking. While this might seem unimportant at first, it convinced me to dig a bit deeper into the underlying structure defining an LLM’s ideology and how it can be expressed through its output. In order to do this, we will also focus on a process called alignment (§3). Building upon multiple works highlighting the political nature of AI and pointing out specific biases, the goal of this research work is to understand the different ideological biases present in LLMs through a deep analysis of the inner workings of LLMs and particularly the alignment process, and to evaluate our findings using a socio-political framework to highlight their societal causes and possible consequences. We are first going to review the previous works on the matter, their approaches to bias analysis and political classification, and try to elaborate new methods to better understand the nature of ideological tendencies in LLMs. We are going to dive into the embedding space and other crucial AI concepts to study how political biases form and how they might be stored, both in the Multi Layer Perceptron (MLP) and the model’s embedding space (Tennenholtz et al., 2024). Then, using both experimental data we gathered for this project and the findings of other papers on the subject, we are going to try and formulate a socio-political hypothesis to explain such biases, the need for and dangers of alignment, and the possible societal impact of political AI.
@@ 3.1 Unsupervised Alignment
As of right now, we don’t possess enough data to precisely measure the profound impact of particular data on a model aligned with unsupervised training since the effects of such training are hard to measure. Some early results we obtained for this work show that such training has an impact on a model’s embedding space, which basically contains the meanings the AI associates with the words we give it. In our tests (Figure 4), an LLM trained on Trotskyist data will have notions like "communism" and "Stalin" farther apart than , for example, a model trained on Marxist-Leninist data  [^4] Along this paper, we will use a lot of alignment references from different types of Marxist-aligned LLMs. The reason for this choice is the abundance of good-quality Marxist data available for free. (mar, ). Training also affects the space in ways we cannot yet understand, changing the position of sometimes completely unrelated notions  [^5] This was written in October 2024, before the first evidence of Emergent Misalignement. See 5.3.
@@ 3.2 ORPO / DPO Alignment
Figure 3: 10101010 most changed words after both unsupervised and ORPO alignment on a Marxist (Trotskyst) dataset. The value used here to measure change is the the distance between base/trained words in the model’s embedding space.
Base Trained
Stalin / Marx 1984 2249
Communism / Authoritarianism 1523 5345
Fascism / Capitalism 2199 385
Figure 4: Norm of the difference vectors of cherry-picked words in a GPT-2 model’s embedding space before and after unsupervized alignment on a Marxist (Trokyst) dataset. Lower distance should mean closer meaning.
This concept of alignment has a lot of similarities with human-directed political influence. Some techniques, such as DPO (Rafailov et al., 2024) or ORPO (Jiwoo Hong and Thorne, 2023), use data sorted into ideological categories and assign weights to them. These techniques aim to modify the AI tendencies to align with them and can evoke quite literally propaganda posters from WWII or the Cold War, mixing desired ideologies with positive signals and the ’enemy’ position with negative brain stimuli. Most big players in today’s LLM market use such techniques to prevent behaviors they find harmful. This method is very powerful as it doesn’t require a lot of training to be effective, but I hypothesize that their impact is more on the ’surface level,’ acting like a role-play prompt rather than completely changing the inner structure of the model.
As for the unsupervised alignment, we also measure a difference in embedding space, which, although smaller, affects a reduced set of vectors in more predictable ways (see Figure 3).
@@ 4.1.2 Absolute Position
In order to obtain more traditional conservative-liberal-Marxist  [^8] We chose these 3 axis because of data availability, and because a simple Left-Right model seemed too simplistic. In the future, it would be interesting to correlate data with actual political entities as done by (Agiza et al., 2024). data on an LLM’s biases, we use a different method that will provide less precise insights but will be grounded in real-world data.
@@ 5 Socio-Political Impacts
In the current state of AI, all major models are produced not by governments but by privately owned corporations. As with social networks, we can hypothesize that this condition will contribute to the power shift from governments to private entities by giving them much more influence. To prevent this, AI regulation would need not only to control an LLM’s output but also precisely measure its biases, and raise awareness around these issues. Indeed, in a lot of countries, political discourse—especially in election times—is heavily regulated (CC, 2012), and as AI can convey ideological messages, it might need to be subjected to the same types of rules.
@@ 5.1 A Marxian Analysis
Our analysis can even be deepened by bringing in the idea of hegemony, developed by Antonio Gramsci (Gramsci, 1971). Indeed, Language Models represent in the current context an essences of the hegemonic ideology of the ruling class. Through the complex tensors of the Multi-layer perceptron, the main points of the capitalist thought framework are encoded, with alignment and/or natural bias. Furthermore, Gramsci theorizes that class hegemony goes through a socio-cultural process that alters ideas themselves. This concept has many similarities with the embedding space tuning in the alignment process: An LLM’s world model is changed not only in its factual aspects but directly in its analysis pathways.  [^10] As our data shows, we can assume that embedding space position data is directly modified by the alignement process, in a predictable way. See 3.1 for our findings.
In a recent paper by (Elsner, 2024), a parallel is drawn between the computational models used in modern AI development and societal superstructures, such as states. Within this framework, Marxian analysis—originally developed to understand and predict the behavior of states and other super-organisms—becomes even more relevant: the alignment problem in AI transforms into a governance problem, closely tied to the interests of the ruling class and the exploitation of the proletariat. This analogy highlights the political stakes of AI alignment and offers insights into potential solutions. By extending the Marxian paradigm beyond mere analysis, we might envision a communist equivalent for computational structures (stateless AI ).
5.2 Kuhn and AI alignement
As a appendix to the Structure of Scientific revolutions, Thomas S. Kuhn tries to go beyond a sociological analysis of the scientific community, and wants to understand the deep pattern ingrained in our brain that lead to the paradigmatic structure of science and the creation of paradigms which are defined not only by subjective opinions and theories, but a completely different understanding and even perception of the world. In order to explain this, Kuhn tries to model the inner working of our brain and more precisely our learning process with a computer program. While this program is nowhere to be found today, he describes it in Second Thoughts as a machine that looks very familiar to someone versed in the technical aspect of modern AI. Indeed, he theorizes a learning process that uses labeled data categories to train a program to perform what we would call today named entity recognition.
Going further than a mere description, he hypothesizes that this simple process is the root cause of the paradigmatic nature of our scientific (and political) knowledge.
5.3 Emergent Misalignement
Since the first version of this paper, a broad range of studies have explored the political alignment problem. While an exhaustive review is beyond the scope of the current work, we will focus on one particular dimension: Emergent Misalignment. As demonstrated earlier through our analytical framework, subtle yet meaningful political biases often manifest along nuances in concepts and ideas that might initially appear entirely unrelated to questions of politics or ethics. Recent studies provide compelling evidence for this hypothesis. In particular, seemingly neutral criteria such as assessments of code quality have been found to correlate strongly with deeper ethical and political divergences (Betley et al., 2025). For a lot of AI researchers, this would suggest the existence of an underlying "evil vector" (aar, ), a latent direction encoding qualities common to all behaviors considered ethically or politically problematic.
Using a Gramscian analysis, we might interpret this alignment vector as encoding fitness relative to a specific cultural hegemony, operating beyond merely political or economic positions and encompassing all dimensions of the prevailing ideological structure. From the perspective of Marxist theory and related critical traditions, this discovery may provide a valuable analytic tool, both for economic and social critique but also elucidating how diverse components of our political system are interconnected and mutually reinforcing.